CHEMILUMINESCENT REACTIONS OF OXYGEN-IODINE SYSTEMS

\underline{JIANDE HAN}, ANATOLY V. KOMISSAROV, SCOTT P. TINNEY AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The reactions of O atoms with iodine containing molecules and the kinetics of the IO radical are of relevance to the removal of tropospheric ozone. These reactions, along with energy transfer processes involving electronically excited states of O_{2}, are also relevant to a new class of oxygen-iodine lasers that employ discharge excitation. Pulsed photolysis of CF_{3}I/N_{2}O/N_{2} mixtures and direct laser excitation of O_{2} in the presence of I_{2} have been used to investigate the iodine oxygen kinetics. The primary findings of this study are that O_{2}(a^{1}Δ) is generated by the reactions IO+O→I+O_{2}(a) and I+I+O_{2}→I_{2}+O_{2}(a). O_{2}(b) and O_{2}(a) are removed by I_{2} with rate constants of 5.8x10^{−11}and < 5x10^{−16} cm^{3} s^{−1}, respectively. The branching fraction for the physical quenching channel O_{2}(b)+I_{2}→O_{2}(a)+I_{2} was found to be 0.4. Re-measurement of the rate constant for O_{2}(a)+O_{2}(X)→2O_{2}(X) yields a value of 9.3x10^{−19} cm^{3} s^{−1}. This result is somewhat lower than previous estimates.